# Temporal Inductive Path Neural Network for Temporal Knowledge Graph Reasoning

### 08 Dec 2023 v2

@@ 6. Conclusion
First, we presented a unified graph, namely the history temporal graph, which represents the comprehensive features of the historical context. This constructed graph allows for more comprehensive utilization of interactions among entities at different historical instances and the temporal feature between historical facts during graph representation learning. Second, we introduced a novel concept of temporal paths, designed to capture query-relevant logical semantic paths on the history temporal graph, which can provide rich structural and temporal context for reasoning tasks. Third, a query-aware temporal path processing framework was also designed, integrated with the introduced temporal edge merging and path aggregation functions. It enables the modeling of temporal path features over the history temporal graph for future temporal fact reasoning.
@@
Overall, the proposed model avoids the need for separate graph learning on each temporal subgraph, making use of the unified graph to represent the information of historical feature to enhance the efficiency of the reasoning process. By starting from the query, capturing and learning over query-aware temporal paths, TiPNN accounts for both structural information and temporal dependencies between entities of separate subgraphs in historical context, and achieves the prediction of future missing facts while offering interpretable reasoning evidence that facilitates users’ analysis of results and model fine-tuning. Besides, in learning historical patterns, the modeling process adopts an entity-independent manner, that means TiPNN doesn’t rely on specific entity representations, enabling it to naturally handle TKG reasoning tasks under the inductive setting.
We have conducted extensive experiments on TKG reasoning benchmark datasets to evaluate the performance of TiPNN. The results demonstrate that the proposed model exhibits superior effectiveness and attains new state-of-the-art achievements.

### 25 Jan 2024 v3

@@ 6 Conclusion
In this work, we have introduced TiPNN, an innovative query-aware temporal path reasoning model for TKGs reasoning tasks, addressing the challenge of predicting future missing temporal facts under the temporal extrapolated setting. First, we presented a unified graph, namely the history temporal graph, which represents the comprehensive features of the historical context. This constructed graph allows for a more comprehensive utilization of temporal features between historical facts during graph representation learning. Second, we introduced a novel concept of temporal paths, designed to capture query-relevant logical semantic paths on the history temporal graph, which can provide rich structural and temporal context for reasoning tasks. Third, a query-aware temporal path processing framework was also designed, integrated with the introduced temporal edge merging and path aggregation functions. It enables the modeling of temporal path features over the history temporal graph for future temporal fact reasoning.
@@
Overall, the proposed model avoids the need for separate graph learning on each temporal subgraph, making use of the unified graph to represent the information of historical feature to enhance the efficiency of the reasoning process. By starting from the query, capturing and learning over query-aware temporal paths, TiPNN accounts for both structural information and temporal dependencies between entities of separate subgraphs in historical context, and achieves the prediction of future missing facts while offering interpretable reasoning evidence that facilitates users’ analysis of results and model fine-tuning. Besides, in learning historical patterns, the modeling process adopts an entity-independent manner, that means TiPNN doesn’t rely on specific entity representations, enabling it to naturally handle TKG reasoning tasks under the inductive setting. We have conducted extensive experiments on TKG reasoning benchmark datasets to evaluate the performance of TiPNN. The results demonstrate that the proposed model exhibits superior effectiveness and attains new state-of-the-art achievements.
The path-based, entity-independent approach employed in TiPNN is well-suited for TKG reasoning tasks as it explores interaction patterns among historical entities, deducing logical semantic interactions that occurred in the past. Although this approach has been applied to traditional KG reasoning  zhu2021neural , the explicit presence of time information and implicit temporal patterns in TKGs renders many excellent KG methods unsuitable for practical applications with a temporal dimension. Moreover, TiPNN consolidates historical subgraph sequences into a unified graph, allowing for more comprehensive utilization of interactions among entities at different times, which helps overcome the limitations and complexities associated with traditional history-evolution methods li2021temporal ; dong2023adaptive when modeling temporal features. While TiPNN provides a novel reference for the TKG reasoning domain, several challenges remain for further research. First, current TKG research is constrained to validation on small-scale datasets, and the performance potential for large-scale datasets is yet to be fully explored. Second, TKGs can be applied to a variety of practical scenarios such as event analysis and trajectory prediction, offering meaningful extensions beyond their current scope. Third, despite the outstanding performance and maturity of existing KG reasoning methods, efficiently integrating advanced KG techniques into the TKG domain poses a significant challenge.
Acknowledgments
This research is funded by the Natural Science Foundation of China under Grant No. 61836013, the Science and Technology Development Fund (FDCT), Macau SAR (file no. 0014/2022/AFJ, 0123/2023/RIA2, 001/2024/SKL), the Start-up Research Grant of University of Macau (File no. SRG2021-00017-IOTSC), the Postdoctoral Fellowship Program of CPSF (No.GZC20232736), the China Postdoctoral Science Foundation Funded Project (No.2023M743565).